Ten quarters dated 1990-1999 are tossed. What is the probability of the following
a) tails occurring on 1991 quarter only (b) tails on all 10 quarters (c) exactly 4 tails (d) at least 3 tails

Answer by stanbon(75887)

Ten quarters dated 1990-1999 are tossed. What is the probability of the following
a) tails occurring on 1991 quarter only
# of ways to succeed: 1
# of possible outcomes 2^10 = 1024
P(tail on 1991 only) = 1/1024
-------------------------------------------
(b) tails on all 10 quarters
(1/2)^10 = 1/1024
------------------------
c) exactly 4 tails
10C4*(1/2)^4*(1/2)^6 = 210/1024
---------------------------------
(d) at least 3 tails
1 - binomcdf(10,0.4,2) = 0.9453
======================================
Cheers,
Stan H.